### v1

@@ 3.2.1 Illustration Example
[…]
@@ 3.3 ATE of True Switchers
As an alternative quantity of interest, we also discuss the ATE of true or actual switching units in the period of the recorded switch or prior. We refer to this quantity as ATE of True Switchers. To this end, recall the unobserved switching variable $S^{\ast}_{i,g,t}$ is one if a person with $D_{i,g,t}^{\ast}=1$ actually switched in that period $t$, and zero otherwise.
Thus, let $N_{S^{*}}:=\sum_{(i,g,t):t\geq 2}\mathbb{I}\left\{S^{*}_{i,g,t}=1\right\}$ denote the number of true switching units. The ATE of true switching units is given by:
@@
As outlined in the introduction, $\delta^{S^{\ast}}$ measures the ATE of all switching-to-treatment units at the time when they actually switch status, across all switching periods among eventually treated groups in the data. This parameter may for instance be of interest when the initial impact of treatment is of particular policy relevance.
We obtain the following result for the identification of the ATE of True Switchers:
@@
Proposition 4 provides a decomposition of the bias of the DID estimand. The first two terms in (9) after $\delta^{S^{\ast}}$ are due to the fact that in the observed switching periods some individuals may in actuality already have undergone treatment in the earlier period and are thus not true switchers. In fact, the third, fourth, and fifth term also arise as there are already switchers in the “pre-treatment” periods prior to the recorded treatment periods. Finally, the last three terms appear because individuals from the comparison groups may actually also be switchers if their treatment occurs in a period just after the comparison.
[…]
Remark 1: Similar to the case of anticipation, our results and methodology are also applicable in setups with spillover effects. For instance, if a policy is enacted at the regional or state level, then units residing in a neighboring geographical location might experience $Y_{i,g,t}(1)$ even though the observed treatment status is recorded as $D_{i,g,t}=0$, and bias results similar to the ones in Propositions 3 and 4 can be derived. As an illustrative example, consider the legalization of recreational marijuana consumption for adults in the U.S., where different states adopted a legalization for recreational usage in different years (e.g., the state of New York in 2021). To the extent that (some) residents from neighboring states without a legalization such as Pennsylvania could “benefit” from the legal purchasing possibilities in New York by crossing the state border, the DID estimator $\widehat{\operatorname{DID}}$ from (7) will exhibit a bias with respect to the ATE of Observed and True Switchers, respectively.
@@ 4 Estimation of ATEs
In this section, we provide modified versions of the $\widehat{\mathrm{DID}}$ estimator to recover both the ATE of Observed Switchers, $\delta^{S}$, and the ATE of True Switchers, $\delta^{S^{\ast}}$. We begin with the former.
@@ 4.1 ATE of Observed Switchers
Similar to the definition of $\widehat{\operatorname{DID}}_{t}$, each of the terms $\widehat{\operatorname{DID}}_{t}^{*}$ and $\widehat{\operatorname{DID}}_{t}^{**}$ is set to zero whenever the corresponding treatment or comparison group does not exist. In particular, we set $\widehat{\operatorname{DID}}_{2}^{*}=0$ and $\widehat{\operatorname{DID}}_{T}^{**}=0$, as no observations are available beyond periods 111 and $T$, respectively.
Heuristically, the estimator $\widehat{\mathrm{DID}}^{S}$ in (10) corrects the original $\widehat{\mathrm{DID}}$ estimator for the two bias terms identified in Proposition 3, using appropriate sample counterparts. Specifically, $\widehat{\operatorname{DID}}_{t}^{*}$ is a “backward-looking correction”: it compares outcome changes from $t-2$ to $t-1$, to adjust for bias due to misclassified treated units in the treatment group at $t-1$. In contrast, $\widehat{\operatorname{DID}}_{t}^{**}$ is a “forward-looking correction”: it compares outcome changes from $t-1$ to $t$ between units still observed as untreated at $t$, but differing in whether they begin treatment at $t+1$, to adjust for bias due to misclassified treated units in the comparison group at $t$. Together, these corrections yield an unbiased estimator of the ATE of observed switchers, $\delta^{S}$.
@@
With staggered adoption designs and the misclassification patterns described in Assumption 7, Assumption 9 ensures that the never treated groups are always correctly classified. If a group indeed begins to receive treatment at $T+1$, it could be misclassified at time $T$, so that this condition is violated. In such cases, the researcher may address the issue by collecting data for one additional period, omitting the final period from analysis, or excluding the affected group (particularly, if the group is not central to the research question and sufficient comparison groups remain). On the other hand, Assumption 10 places a restriction on the data to rule out observed switches in the initial period. Combined with Assumption 7, it ensures that no misclassification occurs at $t=1$, which would otherwise be uncorrectable due to the absence of pre-period information.
@@ 4.2 ATE of True Switchers
Since $\delta^{S^{\ast}}$, the ATE of true switchers, constitutes arguably a more interesting quantity of interest than $\delta^{S}$, we next propose an estimator to recover the second target parameter of interest. To this end, recall the definition of $\delta^{S^{\ast}}$ as:
@@
while $\widehat{\mathrm{DID}}^{S}_{t}$ and $\widehat{\mathrm{DID}}_{t}^{**}$ are given in (10). As before, we adopt the convention that $\widehat{\mathrm{DID}}_{t}^{***}$ is set to zero whenever any of the corresponding groups used therein does not exist in the data. We also set ${\operatorname{DID}}_{1}^{**}=0$, and set $\widehat{\lambda}_{t-1}=0$ when both $\widehat{\operatorname{DID}}_{t-1}^{**}$ and $\widehat{\operatorname{DID}}_{t-1}^{***}$ are zero. Heuristically, the estimator $\widehat{\mathrm{DID}}^{S^{\ast}}$ in (11) aggregates treatment effect estimates from both the observed just-before-switching and switching-to-treatment periods, weighted by the estimated probability of a true switch occurring in each. Specifically, under the restrictions specified below, $\widehat{\mathrm{DID}}^{S^{\dagger}}_{t}$ estimates the unscaled ATE in period $t-1$ for units who actually switched at time $t-1$ but are only observed to switch at time $t$, while $\widehat{\mathrm{DID}}^{S}_{t}$ estimates the unscaled ATE in period $t$ for units whose switch is both actual and observed at time $t$. Moreover, the weight $\widehat{\lambda}_{t-1}$ estimates the share of the former group, and $1-\widehat{\lambda}_{t-1}$ estimates that of the latter.
We require the following assumptions:
@@
Assumption 11 characterizes the behavior of the switching variable $S_{i,g,t}^{\ast}$. It stipulates that $S_{i,g,t}^{\ast}$ depends on the immediate information contained in $D_{g,t}$ and $D_{g,t-1}$, not the entire observed treatment path of group $g$. Moreover, in staggered adoption designs, Assumption 11 complements Assumption 7 by imposing coherent restrictions in terms of the unobserved actual switching indicator $S^{\ast}_{i,g,t}$. Specifically, it rules out the possibility that a true switch, i.e., $S^{*}_{i,g,t}=1$, occurs within a run of treated or untreated periods, except at the first treated period or the last untreated period, respectively. The last part of the assumption parallels Assumption 9 and further rules out such a switch in the last untreated period among the never treated.
The first part of Assumption 12 imposes homogeneity on the switching process across groups and units. It conceptually links the latent treatment status $D^{*}_{i,g,t-1}$ with the switching indicator $S^{*}_{i,g,t-1}$, and requires that all units observed in groups that switch between periods $t-1$ and $t$ share the same probability $\lambda_{t-1}$ of having actually switched in the prior period. [^9] Recall that for notational simplicity, we focus on misclassification occurring only in the period immediately preceding the observed switch. However, the framework can be extended to accommodate more complex settings in which multiple pre-switch periods are affected by misclassification or anticipation. In such cases, one could allow for different pre-treatment misclassification probabilities, though at the cost of more cumbersome notation. The second part formalizes that this total switching probability is fully split between the switching-to-treatment and the just-before-switching periods.
Assumption 13 on the other hand essentially requires that expected unit-level treatment effects cannot be fully heterogeneous, in that they cannot depend on unit’s actual treatment status or switching status. Crucially, this assumption only restricts how such heterogeneity may correlate with actual treatment assignment and switching, while treatment effects may still differ across individuals, groups, or time.
@@
Assumption 14 parallels Assumption 3 from De Chaisemartin and D’Haultefoeuille (2020) and imposes an independence assumption across groups (but not homogeneity). It replaces Assumption 3 and extends it (through the independence) to outcomes in the presence of treatment $Y_{i,g,t}(1)$. Similarly, Assumptions 15 and 16 also represent extensions of Assumptions 4 and 5 to potential outcomes under treatment. They are needed to identify the (unobserved) share of true switchers in a given switching-to-treatment cell, which in turn relies on a comparison with groups that have already switched to treatment in earlier periods. In fact, since standard staggered adoption settings without misspecification do not require such assumptions, these conditions may be viewed as the “price to pay” for not observing the actual switching status. We note however that more general DiD set-ups that also allow for “reverse switchers”, i.e. units switching status back from treatment to no treatment, also rely on Assumptions 14-16 for identification (cf. De Chaisemartin and D’Haultefoeuille, 2020; Fabre, 2023).
@@
Assumption 17 is an asymptotic extension of Assumption 8, that does not only involve comparison groups of the form $D_{g,t}=0$, $D_{g,t-1}=0$, but also all other groups involved in the construction of $\widehat{\mathrm{DID}}^{S^{\ast}}$. Assumption 18 requires that the fourth moments of the potential outcomes $Y_{i,g,t}(d)$ are uniformly bounded across all indexes. It also assumes that the cell sizes $N_{g,t}$ are uniformly bounded. Finally, the assumption ensures the existence of the limiting expectation vector $\lim_{G\to\infty}\mathrm{E}[\mathbf{V}]$ and the asymptotic covariance $\boldsymbol{\Sigma}_{\mathbf{V}}$, both of which are required for establishing the asymptotic distribution of the estimator.
Considering an asymptotic framework where the number of groups $G$ tends to infinity (with the number of time periods $T$ fixed), we have the following asymptotic result:
@@
Under suitable rate conditions, the modified estimator in (13) is a consistent estimator for the ATE of True Switchers for all periods in $\mathcal{T}_{DID,G}$. Alternatively, a continuous, differentiable function could also be used for trimming (e.g., Gutknecht, 2016). Finally, note that the testing procedure outlined in the next section allows to test for the presence of (non-zero) anticipation or misclassification effects, and may thus also inform the researcher whether (12) may be an appropriate assumption or not in a given setting.
@@ 5 Specification Tests
As illustrated in the previous sections, not accounting for misclassification or anticipation in the DID estimator from Equation (7) leads to a bias not only with respect to $\Delta$ in the case of homogeneous treatment effects, but also with respect to both $\delta_{S}$ and $\delta_{S^{\ast}}$, respectively. As a result, a modification of DID is required. However, in the absence of further information, it is not clear how to modify the aforementioned estimator when the researcher does not know how many periods prior to the observed switches are affected specifically. In addition, even if knowledge about affected periods was available, inspection of the proof of Proposition 3 suggests that it is generally difficult to disentangle the bias arising from misclassification or anticipation on the one hand and from a failure of parallel trends (PT) in Assumption 5 on the other. This is, however, important as the implications are different. In particular, while anticipation or misclassification in pre-switch periods can be addressed through a modified DID estimator using earlier periods for comparison (subject to data availability), a failure of strong exogeneity and thus PT is generally invalidating the DID approach for point identification and estimation altogether. [^10] Of course, it may still be possible to adopt a partial identification approach as in Rambachan and Roth (2023) using pre-treatment evidence to bound post-treatment outcome differences. The testing procedure may thus be more informative than standard, event-study type pre-trend tests that do not differentiate between these two different forms of misspecifications.
To address the aforementioned challenges, we propose a two-step testing procedure to detect misspecification due to anticipation effects or misclassifcation, and to guide researchers in the adaptation of the corresponding estimators. In particular, in a first step we aim to maximize confidence in the PT assumption by examining pre-treatment evidence for its validity comparing treated groups with not-yet switched comparison groups in time periods that are known not to be affected by misclassification. Conditional on passing this initial stage, we then conduct a second test for misclassification or anticipation by comparing the same groups using periods that could be misclassified. The outcome of this second test provides guidance on how to adapt the DID estimator to recover $\delta_{S}$ or $\delta_{S^{\ast}}$.
To implement this test procedure, recall that the parallel trends assumption naturally extends to other pre-treatment periods. That is, if for each $(g,k,t)\in\{1,\ldots,G\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$, it holds that
@@
for each $(g,k,t)\in\{1,\ldots,G\}\times\{1,\ldots,G\}\times\{(l+1),\ldots,T\}$ with $l\in\{2,\ldots,T-1\}$. Define the following quantities:
@@
This is a “pre-treatment” statistic comparing pre-treatment periods of switching groups with those same periods of not-yet switching groups. Also, let
$\widehat{\mathrm{\tau}}_{0,0}(t,l)$
This is a “pre-treatment” statistic comparing pre-treatment periods of switching groups with those of not-yet switching groups. To understand the heuristic of the testing procedure, let
@@
We note that in the presence of misspecification (i.e., anticipation or misclassification) and violation of PT, under Assumptions 1-4, 6-8, the expectation of $\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t)$ (provided it exists) is equal to:
$\mathrm{E}\left[\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t)\right]=$ $\mathrm{E}\left[\frac{1}{N_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{g,t}}{N_{g,t-1}}D_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$
@@
On the other hand, for the second statistic, we only have a potential violation of the parallel trends part (by the assumption that misspecification only occurs in cells before the switch):
$\mathrm{E}\left[\widehat{\mathrm{\tau}}_{0,0}^{(l)}(t)\right]=$ $\mathrm{E}\left[\frac{1}{N_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t-2}-Y_{g,t-l}\right)-\frac{1}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t-2}-Y_{g,t-l}\right)\right]$
$=$ $\mathrm{E}\left[\frac{1}{{N}_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\psi_{0,t-2,l}-\frac{1}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\psi_{0,t-2,l}\right].$
This suggests a testing procedure for anticipation or misclassification effects based on moment equalities focusing on the probability limits in the sense of De Chaisemartin and D’Haultefoeuille (2020) or of Callaway and Sant’Anna (2021). Thus, denote:
$\mathrm{\tau}_{1,0}^{(l)}(t)\equiv\lim_{G\rightarrow\infty}\mathrm{E}\left[\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t)\right]$
@@
$\mathrm{\tau}_{0,0}^{(l)}(t)\equiv\lim_{G\rightarrow\infty}\mathrm{E}\left[\widehat{\mathrm{\tau}}_{0,0}^{(l)}(t)\right]$
for every $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$. The testing procedure is based on the following heuristic: if there is pre-treatment evidence suggesting a likely violation of PT, DID should generally not be used for point identification and estimation. On the other hand, if there is pre-treatment evidence supporting PT, we may actually proceed with DID and seek to design the estimator in such a way to avoid misspecification bias. Put differently, we aim to “maximize” our confidence in the PT assumption in a first testing step by looking at pre-treatment evidence in favor of PT, before turning to a misspecification test in a second step to obtain guidance on how to manipulate the estimator to avoid bias. In fact, building on pre-treatment evidence for PT is in line with more recent work by Rambachan and Roth (2023) who exploit pre-treatment PT violations as a way to infer about possible post-treatment violations. Moreover, the procedure is also motivated by the recent findings of De Chaisemartin and D’Haultefoeuille (2024) who show that conducting pre-treatment specification tests may, under the null of correct specification, generally only lead to conservative, but not to too liberal post-treatment inference. This in turn suggest that post-treatment inference carried out conditional on passing pre-treatment specification test, maintains asymptotic size control under the null of correct specification.
To formalize the testing procedure, we first define two distinct tests. The first is for the violation of PT. Specifically, we want to test:
$H_{0}^{PT}:\mathrm{\tau}_{0,0}^{(l)}(t)=0$
@@
$H_{A}^{PT}:\mathrm{\tau}_{0,0}^{(l)}(t)\neq 0$
for some $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$. Similarly, we define the second null hypothesis to be tested as:
$H_{0}^{MC}:\mathrm{\tau}_{1,0}^{(l)}(t)=0$

### v2

@@ 3.2.1 Illustration Example
denotes the observed treatment status of all units other than (i,g,t1)(i,g,t-1) and $(i,g,t)$. Also, for simplicity, we rule out misclassification in any comparison cells for the observed switching-to-treatment cells,
@@ 3.3 ATE of True Switchers
As an alternative estimand, we also consider the ATE for true (or actual) switching units in the period in which treatment is first received. We refer to this parameter as the ATE of true switchers. To define it, recall that the unobserved switching indicator $S^{*}_{i,g,t}$ equals one if unit $i$ in group $g$ truly switches into treatment in period $t$, that is, if $D^{*}_{i,g,t}=1$ and $D^{*}_{i,g,t-1}=0$, and equals zero otherwise.
Let $N_{S^{*}}:=\sum_{(i,g,t):t\geq 2}\mathbb{I}\left\{S^{*}_{i,g,t}=1\right\}$ denote the number of true switching units. The ATE of true switchers is then given by:
@@
As noted in the introduction, $\delta^{S^{*}}$ captures the average treatment effect for units at the time they actually initiate treatment, aggregated across all switching periods among eventually treated groups. This parameter is particularly relevant in applications where the initial impact of treatment is of primary policy interest.
We obtain the following result for identification of the ATE of true switchers.
@@
Proposition 4 provides a detailed decomposition of the bias of the DID estimand relative to the ATE of true switchers. The first two terms following $\delta^{S^{*}}$ arise because, in observed switching periods, some individuals may in fact have already switched into treatment in the preceding period and therefore are not true switchers at time $t$. The subsequent terms reflect analogous contamination in earlier periods that are recorded as pre-treatment, as well as in the comparison groups, where some individuals may actually switch into treatment shortly after the comparison period. Together, these terms illustrate how misalignment between observed and true switching times distorts the DID estimand.
Finally, note that the framework above can equivalently be interpreted as capturing treatment anticipation rather than misclassification. In particular, $Y_{i,g,t-1}(1)$ may represent outcomes realized in period t1t-1 if treatment exposure is anticipated prior to the recorded switch in period $t$. Focusing on anticipation in the period immediately preceding the observed switch for simplicity, Propositions 3 and 4 continue to apply, illustrating that anticipation effects also generate bias in the DID estimand.
Remark 1: In the anticipation interpretation, the setup formally corresponds to a case of “full anticipation,” in the sense that anticipating individuals experience $Y_{i,g,t-1}(1)$ already in period t1t-1. Importantly, however, we place no restriction on the fraction of individuals who anticipate treatment. Moreover, a model with flexible anticipation probabilities and outcomes $Y_{i,g,t-1}(1)$ is observationally equivalent to an alternative model featuring “fractional anticipation” combined with different anticipation probabilities. For this reason, we do not view the full-anticipation formulation as restrictive.
Remark 2: Similar to the case of anticipation, our results and methodology are also applicable in setups with spillover effects. For example, when a policy is implemented at the regional or state level, units in neighboring locations may experience treated potential outcomes $Y_{i,g,t}(1)$ even though their observed treatment status remains $D_{i,g,t}=0$. In such cases, bias expressions analogous to those in Propositions 3 and 4 can be derived. As an illustrative example, consider the legalization of recreational marijuana consumption for adults in the U.S., where different states adopted legalization in different years, such as New York in 2021. To the extent that some residents of neighboring non-legalizing states, such as Pennsylvania, can benefit from legal purchasing opportunities by crossing state borders, the DID estimator $\widehat{\mathrm{DID}}$ in (7) will generally be biased relative to the ATE of observed switchers and the ATE of true switchers.
@@ 4 Estimation of ATEs
In this section, we propose modified versions of the $\widehat{\mathrm{DID}}$ estimator that recover both the ATE of observed switchers, $\delta^{S}$, and the ATE of true switchers, $\delta^{S^{*}}$. We begin with estimation of the former.
@@ 4.1 ATE of Observed Switchers
As with $\widehat{\mathrm{DID}}_{t}$, each of the correction terms $\widehat{\mathrm{DID}}^{*}_{t}$ and $\widehat{\mathrm{DID}}^{**}_{t}$ is set to zero whenever the corresponding treatment or comparison group does not exist. In particular, we set $\widehat{\mathrm{DID}}^{*}_{2}=0$ and $\widehat{\mathrm{DID}}^{**}_{T}=0$, since observations prior to period $1$ and after period $T$ are unavailable.
Heuristically, the estimator $\widehat{\mathrm{DID}}^{S}$ in (10) corrects the original $\widehat{\mathrm{DID}}$ estimator for the two sources of bias identified in Proposition 3, using appropriate sample counterparts. Specifically, $\widehat{\operatorname{DID}}_{t}^{*}$ is a “backward-looking correction”: it compares outcome changes from t2t-2 to t1t-1, to adjust for bias due to misclassified treated units in the treatment group at t1t-1. In contrast, $\widehat{\operatorname{DID}}_{t}^{**}$ is a “forward-looking correction”: it compares outcome changes from t1t-1 to $t$ between units still observed as untreated at $t$, but differing in whether they begin treatment at $t+1$, to adjust for bias due to misclassified treated units in the comparison group at $t$. Together, these corrections yield an unbiased estimator of the ATE of observed switchers, $\delta^{S}$.
@@
In staggered adoption designs with the misclassification patterns misclassification patterns described in Assumption 7, Assumption 9 ensures that the never treated groups are always correctly classified. If a group indeed begins to receive treatment at $T+1$, it could be misclassified at time $T$, so that this condition is violated. In such cases, the researcher may address the issue by collecting data for one additional period, omitting the final period from analysis, or excluding the affected group (particularly, if the group is not central to the research question and sufficient comparison groups remain). On the other hand, Assumption 10 places a restriction on the data to rule out observed switches in the initial period. Combined with Assumption 7, it ensures that no misclassification occurs at $t=1$, which would otherwise be uncorrectable due to the absence of pre-period information. Assumption 10 rules out observed treatment switches in the initial period. When combined with Assumption 7, it guarantees that no misclassification occurs at $t=1$, which would otherwise be uncorrectable due to the absence of pre-treatment information.
@@
Remark 3: Throughout the paper, we ignore the presence of additional observed control variables $\mathbf{X}_{i,gt}$. In fact, when $\mathbf{X}_{i,gt}$ is discrete with a finite number of support points, we can readily modify all assumptions to hold also conditional on $\mathbf{X}_{i,gt}$. Together with a suitable overlap condition (Khan2010), we may then construct the estimator $\widehat{\mathrm{DID}}^{S}$ for each value of the covariate vector to recover an estimate of the ATE of Observed Switchers conditional $\mathbf{X}_{i,11}$. Alternatively, we may also aggregate the estimates from each cell to obtain an estimate of the unconditional ATE of Observed Switchers $\delta^{S}$. When $\mathbf{X}_{i,gt}$ is continuous or discrete with infinite support, estimation is more involved (e.g., Abadie2005).
@@ 4.2 ATE of True Switchers
Since $\delta^{S^{*}}$, the ATE of true switchers, is arguably a more substantively meaningful estimand than $\delta^{S}$, we now propose an estimator for this second target parameter. Recall that $\delta^{S^{*}}$ is defined as
@@
while $\widehat{\mathrm{DID}}^{S}_{t}$ and $\widehat{\mathrm{DID}}_{t}^{**}$ are given in (10). As before, we adopt the convention that $\widehat{\mathrm{DID}}^{***}_{t}$ is set to zero whenever any of the corresponding groups do not exist in the data. We also set $\widehat{\mathrm{DID}}^{**}_{1}=0$, and define $\widehat{\lambda}_{t-1}=0$ whenever both $\widehat{\mathrm{DID}}^{**}_{t-1}$ and $\widehat{\mathrm{DID}}^{***}_{t-1}$ equal zero. Heuristically, the estimator $\widehat{\mathrm{DID}}^{S^{\ast}}$ in (11) aggregates treatment effect estimates from both the observed just-before-switching and switching-to-treatment periods, weighted by the estimated probability of a true switch occurring in each. Under the identifying restrictions introduced below, $\widehat{\mathrm{DID}}^{S^{\dagger}}_{t-1}$ estimates the (unscaled) ATE in period t1t-1 for units that truly switch in t1t-1 but are only observed to switch in $t$, whereas $\widehat{\mathrm{DID}}^{S}_{t}$ estimates the (unscaled) ATE in period $t$ for units whose switch is both actual and observed at time $t$. Moreover, the weight $\widehat{\lambda}_{t-1}$ estimates the share of the former group, while $1-\widehat{\lambda}_{t-1}$ estimates the share of the latter.
We impose the following additional assumptions.
@@
Assumption 11 governs the behavior of the latent switching indicator $S^{*}_{i,g,t}$. It restricts true switching to be determined solely by the local information contained in $(D_{g,t-1},D_{g,t})$, rather than by the full observed treatment path of group $g$. In staggered adoption designs, Assumption 11 complements Assumption 7 by imposing coherent restrictions in terms of the unobserved actual switching indicator $S^{\ast}_{i,g,t}$. Specifically, it rules out the possibility that a true switch, i.e., $S^{*}_{i,g,t}=1$, occurs within a run of treated or untreated periods, except at the first treated period or the last untreated period, respectively. The last part of the assumption parallels Assumption 9 and further rules out such a switch in the last untreated period among the never treated.
Assumption 12 imposes homogeneity on the switching process. The first part links the latent treatment status $D^{*}_{i,g,t-1}$ to the switching indicator $S^{*}_{i,g,t-1}$ and requires that, among groups observed to switch between periods t1t-1 and $t$, all units share the same probability $\lambda_{t-1}$ of having actually switched in the earlier period. [^7] Recall that for notational simplicity, we restrict attention to misclassification or anticipation occurring only in the period immediately preceding the observed switch. The framework can be extended to allow for multiple pre-switch periods with distinct switching probabilities, at the cost of more involved notation. The second part ensures that the total switching probability is fully allocated between the just-before-switching period and the observed switching period.
Finally, Assumption 13 restricts treatment effect heterogeneity. It requires that expected unit-level treatment effects do not depend on a unit’s actual treatment status or switching status, although they may still vary across individuals, groups, or time. Thus, the assumption limits only how treatment effect heterogeneity may correlate with latent treatment assignment and switching, rather than ruling out heterogeneity altogether.
@@
Assumption 14 parallels Assumption 3 from CH2020 and imposes an independence assumption across groups (but not homogeneity). It replaces Assumption 3 and extends it (through the independence) to outcomes in the presence of treatment $Y_{i,g,t}(1)$. Similarly, Assumptions 15 and 16 also represent extensions of Assumptions 4 and 5 to potential outcomes under treatment. They are needed to identify the (unobserved) share of true switchers in a given switching-to-treatment cell, which in turn relies on a comparison with groups that have already switched to treatment in earlier periods. In fact, since standard staggered adoption settings without misspecification do not require such assumptions, these conditions may be viewed as the “price to pay” for not observing the actual switching status. We note however that more general DiD set-ups that also allow for “reverse switchers”, i.e. units switching status back from treatment to no treatment, also rely on Assumptions 14-16 for identification (cf. CH2020; F2023).
@@
Assumption 17 is an asymptotic analogue of Assumption 8 and ensures the existence of all comparison groups required to construct $\widehat{\mathrm{DID}}^{S^{\ast}}$. Assumption 18 requires that the fourth moments of the potential outcomes $Y_{i,g,t}(d)$ are uniformly bounded across all indexes. It also assumes that the cell sizes $N_{g,t}$ are uniformly bounded. Finally, the assumption ensures the existence of the limiting expectation vector $\lim_{G\to\infty}\mathrm{E}[\mathbf{V}]$ and the asymptotic covariance $\boldsymbol{\Sigma}_{\mathbf{V}}$, both of which are required for establishing the asymptotic distribution of the estimator.
Considering an asymptotic framework in which the number of groups $G$ tends to infinity while the number of time periods $T$ remains fixed, we obtain the following result.
@@
Under suitable rate conditions, the modified estimator in (13) is a consistent estimator for the ATE of true switchers for all periods in $\mathcal{T}_{DID,G}$. Alternatively, a continuous, differentiable function could also be used for trimming (e.g., G2016). Finally, note that the testing procedure outlined in the next section allows to test for the presence of (non-zero) anticipation or misclassification effects, and may thus also inform the researcher about the suitability of (12).
@@ 5 Specification Tests
As illustrated in the previous sections, not accounting for misclassification or anticipation in the DID estimator from Equation (7) leads to a bias not only with respect to $\Delta$ in the case of homogeneous treatment effects, but also with respect to both $\delta_{S}$ and $\delta_{S^{\ast}}$, respectively. As a result, a modification of DID is required. However, in the absence of further information, it is not clear how to modify the aforementioned estimator when the researcher does not know how many periods prior to the observed switches are affected. In addition, even if knowledge about affected periods was available, inspection of the proof of Proposition 3 suggests that it is generally difficult to disentangle the bias arising from misclassification or anticipation on the one hand and from a failure of parallel trends (PT) in Assumption 5 on the other. This is, however, important as the implications are different. In particular, while anticipation or misclassification in pre-switch periods can be addressed through a modified DID estimator using earlier periods for comparison (subject to data availability), a failure of strong exogeneity and thus PT generally invalidates the DID approach for point identification and estimation altogether. [^8] Of course, it may still be possible to adopt a partial identification approach as in Rambachan2023 using pre-treatment evidence to bound post-treatment outcome differences.
To address these challenges, we propose two specification tests based on moment equalities to detect violations of PT in pre-treatment periods on the one hand, and misspecification due to anticipation effects or misclassifcation on the other. In particular, the first test compares outcome averages from switching groups with not-yet switched groups in periods prior to a threshold date assumed to be known and not affected by misclassification. [^9] For instance, in the context of the study of Bindler2018, who analyzed the effects of the (staggered) abolition of capital punishment across different offense categories on jury severity, one may use periods that preceed the enactment of these law changes. The second test is conceptually identical, but instead focuses on comparisons from periods potentially affected by misclassification. Heuristically, a researcher may therefore use the first test to gather evidence for pre-treatment violations of PT, and then conduct the second test for misclassification comparing the same groups using periods that could (potentially) be misclassified. Moreover, as both test rely on moment comparisons and can exploit all possible combinations from the time periods considered, they are more likely to detect violations of PT or misclassification in pre-treatment periods than conventional event-study based specification tests which just check for non-zero effects relative to a given reference period (see Remark 6 below).
Formally, recall that the parallel trends assumption naturally extends to other pre-treatment periods. That is, if for each $(g,k,t)\in\{1,\ldots,G\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$:
@@
for each $(g,k,t)\in\{1,\ldots,G\}\times\{1,\ldots,G\}\times\{(l+1),\ldots,T\}$ with $l\in\{2,\ldots,T-1\}$. We may use this insight to define the following empirical moment comparisons:
$\widehat{\mathrm{\tau}}_{0,0}(t,l)$
and
@@
While (14) is a pre-treatment statistic that compares outcomes from switching groups with those of not-yet switched groups for periods that pre-date potential misclassification in t1t-1, namely t2t-2 and tlt-l with $l>2$, (15) is a pre-treatment statistic that compares those outcomes using period t1t-1 as reference period. To understand the heuristic, note that because of the assumption that misspecification only occurs in t1t-1, under Assumptions 1-4, 6-8 we have a potential violation of PT as the only “bias source”:
$\mathrm{E}\left[\widehat{\mathrm{\tau}}_{0,0}(t,l)\right]=$ $\mathrm{E}\left[\frac{1}{N_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t-2}-Y_{g,t-l}\right)-\frac{1}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t-2}-Y_{g,t-l}\right)\right]$
$=$ $\mathrm{E}\left[\frac{1}{{N}_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\psi_{0,t-2,l}-\frac{1}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\psi_{0,t-2,l}\right],$
with
@@
On the other hand, the expectation of $\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t)$ (provided it exists) is equal to:
$\mathrm{E}\left[\widehat{\mathrm{\tau}}_{1,0}(1,l)\right]=$ $\mathrm{E}\left[\frac{1}{N_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{g,t}}{N_{g,t-1}}D_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$
@@
This suggests basing the tests for pre-treatment PT and misclassification on moment equalities. Thus, define:
$\mathrm{\tau}_{0,0}(t,l)\equiv\lim_{G\rightarrow\infty}\mathrm{E}\left[\widehat{\mathrm{\tau}}_{0,0}(t,l)\right]$
@@
$\mathrm{\tau}_{1,0}(t,l)\equiv\lim_{G\rightarrow\infty}\mathrm{E}\left[\widehat{\mathrm{\tau}}_{1,0}(t,l)\right]$
for every $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$. Specifically, to test for violations of PT, we want to test that:
$H_{0}^{PT}:\mathrm{\tau}_{0,0}(t,l)=0$
@@
$H_{A}^{PT}:\mathrm{\tau}_{0,0}(t,l)\neq 0$
for some $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$. Similarly, the second null hypothesis about the presence of misclassification or anticipation is given by:
$H_{0}^{MC}:\mathrm{\tau}_{1,0}(t,l)=0$